In a sample of 1500 new iPhones, 12 were found to be defective. At this rate, how many defective iPhones would be expected in a sample of 100,000 new iPhones? Use a proportion to set up and solve this problem.
Answer by FreezyCold(1)

12 x
-- = --
1500 100000
cross multiple
12*100000=1500*x
-----------------
1200000=1500x
-------------
inverse operation
---------------
1200000/1500=x
--------------
800=x